OpenStudy (anonymous):

Find the value(s) of α such that a = 2i-j+αk is three times as long as b = 2i+j-k.

OpenStudy (anonymous):

\[|a|=\sqrt{4+1+\alpha} =3\sqrt{4+1+1}=|b|\] 4+1+alpha = 9(4+1+1) alpha = 36 +1+1 -4-1 =33

OpenStudy (anonymous):

Shouldn't it be α^2 within |a| since you squared all the values of vectors a and b within the square root.